# Designing Phase Sensitive Probes of Monopole Superconducting Order

### v1

@@ II First order Josephson coupling in the presence of spin-orbit coupling
where $T^{(b)}_{0}(\mathbf{k},\mathbf{k}^{\prime})=T_{0}(\mathbf{k},\mathbf{k}^{\prime})(u^{*}_{R,0}(\mathbf{k})u_{L,0}(\mathbf{k}^{\prime})+\mathbf{u}^{*}_{R}(\mathbf{k})\cdot\mathbf{u}_{L}(\mathbf{k}^{\prime}))$ and $\mathbf{T}^{(b)}(\mathbf{k},\mathbf{k}^{\prime})=T_{0}(\mathbf{k},\mathbf{k}^{\prime})(u^{*}_{R,0}(\mathbf{k})\mathbf{u}_{L}(\mathbf{k}^{\prime})+\mathbf{u}^{*}_{R}(\mathbf{k}){u}_{L,0}(\mathbf{k}^{\prime})+i\mathbf{u}^{*}_{R}(\mathbf{k})\times\mathbf{u}_{L}(\mathbf{k}^{\prime}))$. For most parts of the manuscript, we consider the case of zero bare interface spin-orbit interaction for simplicity. We discuss the results when the bare interface spin-orbit coupling $\mathbf{T}(\mathbf{k},\mathbf{k}^{\prime})\neq 0$ in Sec. VI.
Furthermore, in a weak-coupling superconductor, we can project its pairing order to low energy states near the Fermi surfaces participating in Cooper pairing. For pairing order already expressed in the band representation $\Delta_{\alpha}^{(b)}(\mathbf{k})=U_{\alpha}^{\dagger}(\mathbf{k})\Delta_{\alpha}(\mathbf{k})U^{*}_{\alpha}(-\mathbf{k})$, we capture the key ingredients of the low energy pairing by focusing on the Fermi surface projected pairing order in this representation, defining $\Delta^{(bp)}_{\alpha}(\mathbf{k})\equiv P_{+}^{(b)}\Delta_{\alpha}^{(b)}(\mathbf{k})P_{+}^{(b)}$, where $P_{+}^{(b)}=\mathrm{diag}(1,0)$ is the projection operator in band representation. Here, without loss of generality, we have ordered the band basis so that the states at the Fermi surface of interest correspond to the first eigenvalues in the particle and hole parts (see Appendix B). The superscript “$bp$” denotes the band diagonal basis with bands projected to the Fermi surfaces of interest. The anomalous Green’s function of the corresponding BdG Hamiltonian in this band projected basis takes a simple form, $\mathcal{F}^{(bp)}_{\alpha}(\mathbf{k},ip_{n})\approx-e^{i\phi_{\alpha}}\Delta_{\alpha}^{(bp)}(\mathbf{k})/(p_{n}^{2}+E^{(bp)2}_{\alpha,\mathbf{k}})$. This form is analogous to the simplified form of $\mathcal{F}_{\alpha}(\mathbf{k},ip_{n})$ in the case of spin degenerate band and unitary pairing. However, here, the energy of the BdG quasiparticle is given by $E^{(bp)}_{\alpha,\mathbf{k}}=\sqrt{\xi_{\alpha,\mathbf{k}}^{(bp)2}+\mathrm{Tr}(\Delta^{(bp)\dagger}_{\alpha}(\mathbf{k})\Delta^{(bp)}_{\alpha}(\mathbf{k}))}$, where $\xi^{(bp)}_{\alpha,\mathbf{k}}$ is the dispersion of the band that forms the Fermi surface. The expression of the form factor in Eq. (8) now reduces to $\mathfrak{F}(\mathbf{k},\mathbf{k}^{\prime})=\mathrm{Tr}[\Delta_{L}^{(bp)}(\mathbf{k}^{\prime})T^{(b)\mathrm{T}}(-\mathbf{k},-\mathbf{k}^{\prime})\Delta_{R}^{(bp)*}(\mathbf{k})T^{(b)}(\mathbf{k},\mathbf{k}^{\prime})]$, where the pairing and tunneling matrices are in the projected band basis. We use this band projected form factor to analyze Josephson junctions of monopole superconductors, as demonstrated in the following section.
@@ III Josephson coupling in junctions of monopole superconductors
Having discussed the form of the first order Josephson current in a general SIS junction, our focus turns to junctions that include monopole superconductors. This discussion unfolds in two different approaches. We first derive the form factor that determines the Josephson current using the microscopic derivation based on linear response theory, as discussed in Sec. II. We work in the weak-coupling regime, projecting the pairing order to low energy states that comprise the Fermi surface. Secondly, we supplement the first approach with a discussion based on the underlying symmetry and topological principles, which are essential in designing a set of Josephson junctions to probe the monopole superconducting order.
@@ III.1 Microscopic derivation of Josephson coupling for monopole superconductors
We first extend the results from linear response theory in Sec. II to determine the first order Josephson current for junctions that include monopole superconductors. Monopole pairing order is an exotic class of topologically obstructed pairing order that arises from the nontrivial Berry phase of Cooper pairs. This, for example, can arise when pairing occurs between two Fermi surfaces $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$ with different Chern numbers, $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ respectively, in doped time-reversal symmetry broken Weyl semimetals in three dimensions [1]. In the weak-coupling regime, we describe this system using mean-field BdG Hamiltonian in Eq. (2), with its Nambu spinor modified to $\Psi(\mathbf{k})=(c_{1;\mathbf{k},\uparrow},c_{1;\mathbf{k},\downarrow},c_{2;-\mathbf{k},\uparrow}^{\dagger},c_{2;-\mathbf{k},\downarrow}^{\dagger})^{\mathrm{T}}$ to describe the inter-Fermi surface pairing $\Delta_{\mathrm{inter}}(\mathbf{k})$ between parity-related $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$, where $c_{i;\mathbf{k},\sigma}$ denotes the annihilation operator of an electron on $\mathrm{FS}_{i}$ ($i=1,2$) with momentum $\mathbf{k}$ and spin $\sigma=\uparrow,\downarrow$. Correspondingly, the BdG Hamiltonian kernel takes the following form,
@@
The inter-Fermi surface pairing, $\Delta_{\mathrm{inter};\sigma,\sigma^{\prime}}(\mathbf{k})\sim\langle c_{2;-\mathbf{k},\sigma^{\prime}}c_{1;\mathbf{k},\sigma}\rangle$ in the spin-$\uparrow$,$\downarrow$ , ↓ representation, describes the zero center-of-mass momentum pairing, since FS1 and FS2 are related by parity. In the weak-coupling regime, superconducting pairing occurs at low energy close to the Fermi surfaces; hence, we express the pairing in the helical band representation as $\Delta_{\mathrm{inter};s,s^{\prime}}^{(b)}(\tilde{\mathbf{k}})\sim\langle\psi_{2;-\mathbf{\tilde{k}},s^{\prime}}\psi_{1;\mathbf{\tilde{k}},s}\rangle$, where $\psi_{1;\mathbf{\tilde{k}},s}=\sum_{\sigma=\uparrow,\downarrow}\chi_{1;s,\sigma}^{*}(\tilde{\mathbf{k}})c_{1;\tilde{\mathbf{k}}+\mathbf{K}_{0},\sigma}$ and $\psi_{2;\mathbf{\tilde{k}},s}=\sum_{\sigma=\uparrow,\downarrow}\chi_{2;s,\sigma}^{*}(\tilde{\mathbf{k}})c_{2;\tilde{\mathbf{k}}-\mathbf{K}_{0},\sigma}$ are annihilation operators of helical band eigenstates of $\mathcal{H}_{\mathrm{kin},1}$ and $\mathcal{H}_{\mathrm{kin},2}$ respectively. When the Chern numbers of two Fermi surfaces are different, the inter-Fermi surface pairing inherits the band topology of single-particle states at $\mathbf{k}$ and $-\mathbf{k}$ in a nontrivial way, and the Cooper pair scatters in the presence of a magnetic monopole in momentum space given by the pair monopole charge, $q_{p}=(\mathcal{C}_{1}-\mathcal{C}_{2})/2$. Consequently, this leads to a topologically obstructed superconducting pairing order that cannot be well defined over an entire Fermi surface. Instead, the pairing order is described by monopole harmonics [11, 37] that all belong to the topological sector characterized by the pair monopole charge $q_{p}$ [1].
Upon projecting to the helical Fermi surfaces in the weak-coupling limit, the effective pairing in the helical band-diagonal Nambu basis $(\psi_{1;\tilde{\mathbf{k}},+},\psi^{\dagger}_{2;-\tilde{\mathbf{k}},+})^{T}$ is
@@
$=\sqrt{2\pi}\left(\mathcal{Y}_{\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}}),-\mathcal{Y}_{\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\right)^{T},$
and is related to states on $\mathrm{FS}_{2}$ by $\chi_{2,+}(\tilde{\mathbf{k}})=\sigma_{z}\chi_{1,+}(-\tilde{\mathbf{k}})$. Here, $\bm{\Omega}_{\tilde{{k}}}=(\theta_{\tilde{k}},\varphi_{\tilde{k}})$, with $\theta_{\tilde{k}}$ and $\varphi_{\tilde{k}}$ being the polar and azimuthal angles of the spherical Fermi surfaces. $\mathcal{Y}_{q;l,l_{z}}(\bm{\Omega}_{\tilde{k}})$ is the monopole harmonic function [11, 37] (reviewed in Appendix C), now defined in momentum space. The monopole harmonic function is labelled by the monopole charge $q$, the eigenvalue of angular momentum $l$, and the eigenvalue of $z$-component angular momentum $l_{z}$. Above, $\lambda=\pm 1$ corresponds to two different choices of gauge. The upper (lower) sign denotes the choice in which the state $\chi_{1,+}(\tilde{\mathbf{k}})$ is well-defined locally near the north (south) pole of $\mathrm{FS}_{1}$, while $\chi_{2,+}(\tilde{\mathbf{k}})$ is well-defined near the south (north) pole of $\mathrm{FS}_{2}$, with the Dirac strings located at the respective antipodal points of these two Fermi surfaces. The Chern numbers of $\chi_{1,+}$ on $\mathrm{FS}_{1}$ and $\chi_{2,+}$ on $\mathrm{FS}_{2}$ are $\mathcal{C}_{1}=-1$ and $\mathcal{C}_{2}=+1$ respectively. The unitary transformations that diagonalize the Weyl band Hamiltonians $\mathcal{H}_{\mathrm{kin}}^{(\nu=\pm 1)}$ in Eq. (19) take the form
@@
in which $\mathcal{Y}^{*}_{q;l,l_{z}}=(-1)^{q+l_{z}}\mathcal{Y}_{-q;l,-l_{z}}$.
@@
We now derive the form factor in the first order Josephson coupling between a spin singlet superconductor with a spin-independent band Hamiltonian at the left side and the $q_{p}=-1$ monopole superconductor at the right side, in which there is only bare spin-independent tunneling at the barrier. In the band-projected basis, the monopole pairing order is described by an effective $d$-vector shown in Eq. (17). The form factor hence takes the form
@@
$T^{(b)}_{0}(\pm\mathbf{k},\pm\mathbf{k}^{\prime})$ $=T_{0}(\pm\mathbf{k},\pm\mathbf{k}^{\prime})[u_{R,0}^{(\pm\mathbf{K}_{0})}(\pm\tilde{\mathbf{k}})]^{*}$ (26a)
@@
Above, $(u_{R,0}^{(\pm\mathbf{K}_{0})}(\tilde{\mathbf{k}}),\mathbf{u}_{R}^{(\pm\mathbf{K}_{0})}(\tilde{\mathbf{k}}))$ are the complex four-vectors defined in Eq. (22).
Consequently, the form factor describing the first order Josephson coupling between a $q_{p}=-1$ monopole superconductor and a singlet superconductor simplifies to
@@
×$\times{T_{0}(-\mathbf{k},-\mathbf{k}^{\prime})}{T_{0}(\mathbf{k},\mathbf{k}^{\prime})}\mathcal{Y}_{-\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\mathcal{Y}_{-\frac{1}{2};\frac{1}{2},\frac{1}{2}}(\bm{\Omega}_{\tilde{k}}),$
in which $\Delta_{L,0}(\mathbf{k}^{\prime})=d_{L,0}(\mathbf{k}^{\prime})$, $\Delta_{R,\mathrm{MSC}}^{{(q_{p}=-1)}}(\tilde{\mathbf{k}})$ is the band-projected effective monopole harmonic pairing order in Eq. (23) characterized by the pair monopole charge $q_{p}=-1$, and $\tilde{\mathbf{k}}=\mathbf{k}-\mathbf{K}_{0}$ is the wave vector with respect to the Weyl node. Despite the form factor expressed in Eq. (27) including multiple monopole harmonic functions, $\mathfrak{F}_{\mathrm{sing}-\mathrm{MSC}}$ is always non-singular, as it is related to the Josephson current, a physical observable. By employing the addition theorem of monopole harmonic functions [37] (see also Appendix C), which states that the product of two monopole harmonic functions with charges $q_{1}$ and $q_{2}$ can be expressed as a sum of monopole harmonic functions with monopole charge $q_{3}=q_{1}+q_{2}$, we find that the form factor in Eq. (27) has zero total monopole charge. In other words, it can be represented using standard spherical harmonic functions without singularities over the Fermi surface.
@@
$\mathfrak{F}_{\mathrm{trip}-\mathrm{MSC}}^{(q_{p}=-1)}(\tilde{\mathbf{k}},\mathbf{k}^{\prime})=-2\pi{T_{0}(-\mathbf{k},-\mathbf{k}^{\prime})}{T_{0}(\mathbf{k},\mathbf{k}^{\prime})}\times$
${\Delta_{R,\mathrm{MSC}}^{{(q_{p}=-1)}^{*}}(\tilde{\mathbf{k}})}[(d_{L,x}(\mathbf{k}^{\prime})-id_{L,y}(\mathbf{k}^{\prime}))\mathcal{Y}_{-\frac{1}{2};\frac{1}{2},-\frac{1}{2}}^{2}(\bm{\Omega}_{\tilde{k}})$
@@
It depends on, $\mathbf{d}_{L}(\mathbf{k}^{\prime})$, the $d$-vector which encodes the spin structure of the triplet pairing order at the left. In contrast, the form factor $\mathfrak{F}_{\mathrm{sing}-\mathrm{MSC}}^{(q_{p}=-1)}$ in Eq. (27) takes a simpler form, as the symmetry of the singlet pairing order at the left is primarily determined by its symmetry in momentum space. For the microscopic models studied later in Sec. IV and VI, to identify the momentum space topology of the monopole pairing order, we will focus on junctions with a spin singlet superconductor for simplicity.

### v2

@@ II First-order Josephson coupling in the presence of spin-orbit coupling
where $T^{(b)}_{0}(\mathbf{k},\mathbf{k}^{\prime})=T_{0}(\mathbf{k},\mathbf{k}^{\prime})(u^{*}_{R,0}(\mathbf{k})u_{L,0}(\mathbf{k}^{\prime})+\mathbf{u}^{*}_{R}(\mathbf{k})\cdot\mathbf{u}_{L}(\mathbf{k}^{\prime}))$ and $\mathbf{T}^{(b)}(\mathbf{k},\mathbf{k}^{\prime})=T_{0}(\mathbf{k},\mathbf{k}^{\prime})(u^{*}_{R,0}(\mathbf{k})\mathbf{u}_{L}(\mathbf{k}^{\prime})+\mathbf{u}^{*}_{R}(\mathbf{k}){u}_{L,0}(\mathbf{k}^{\prime})+i\mathbf{u}^{*}_{R}(\mathbf{k})\times\mathbf{u}_{L}(\mathbf{k}^{\prime}))$. For most parts of the manuscript, we consider the case of zero bare interface spin-orbit interaction for simplicity. We discuss the results when the bare interface spin-orbit coupling $\mathbf{T}(\mathbf{k},\mathbf{k}^{\prime})\neq 0$ in Sec. VII.
Furthermore, in a weak-coupling superconductor, we can project its pairing order to low-energy states near the Fermi surfaces participating in Cooper pairing. For pairing order already expressed in the band representation $\Delta_{\alpha}^{(b)}(\mathbf{k})=U_{\alpha}^{\dagger}(\mathbf{k})\Delta_{\alpha}(\mathbf{k})U^{*}_{\alpha}(-\mathbf{k})$, we capture the key ingredients of the low-energy pairing by focusing on the Fermi surface projected pairing order in this representation, defining $\Delta^{(bp)}_{\alpha}(\mathbf{k})\equiv P_{+}^{(b)}\Delta_{\alpha}^{(b)}(\mathbf{k})P_{+}^{(b)}$, where $P_{+}^{(b)}=\mathrm{diag}(1,0)$ is the projection operator in band representation. Here, without loss of generality, we have ordered the band basis so that the states at the Fermi surface of interest correspond to the first eigenvalues in the particle and hole parts (see Appendix B). The superscript “$bp$” denotes the band diagonal basis with bands projected to the Fermi surfaces of interest. The anomalous Green’s function of the corresponding BdG Hamiltonian in this band projected basis takes a simple form, $\mathcal{F}^{(bp)}_{\alpha}(\mathbf{k},ip_{n})\approx-e^{i\phi_{\alpha}}\Delta_{\alpha}^{(bp)}(\mathbf{k})/(p_{n}^{2}+E^{(bp)2}_{\alpha,\mathbf{k}})$. This form is analogous to the simplified form of $\mathcal{F}_{\alpha}(\mathbf{k},ip_{n})$ in the case of spin degenerate band and unitary pairing. However, here, the energy of the BdG quasiparticle is given by $E^{(bp)}_{\alpha,\mathbf{k}}=\sqrt{\xi_{\alpha,\mathbf{k}}^{(bp)2}+\mathrm{Tr}(\Delta^{(bp)\dagger}_{\alpha}(\mathbf{k})\Delta^{(bp)}_{\alpha}(\mathbf{k}))}$, where $\xi^{(bp)}_{\alpha,\mathbf{k}}$ is the dispersion of the band that forms the Fermi surface. The expression of the form factor in Eq. (8) now reduces to $\mathfrak{F}(\mathbf{k},\mathbf{k}^{\prime})=\mathrm{Tr}[\Delta_{L}^{(bp)}(\mathbf{k}^{\prime})T^{(b)\mathrm{T}}(-\mathbf{k},-\mathbf{k}^{\prime})\Delta_{R}^{(bp)*}(\mathbf{k})T^{(b)}(\mathbf{k},\mathbf{k}^{\prime})]$, where the pairing and tunneling matrices are in the projected band basis. We use this band projected form factor to analyze Josephson junctions of monopole superconductors, as demonstrated in the following section.
@@ III Josephson coupling in junctions of monopole superconductors
Having discussed the form of the first-order Josephson current in a general SIS junction, our focus turns to junctions that include monopole superconductors. This discussion unfolds in two different approaches. We first derive the form factor that determines the Josephson current using the microscopic derivation based on linear response theory, as discussed in Sec. II. We work in the weak-coupling regime, projecting the pairing order to low-energy states that comprise the Fermi surface. Secondly, we supplement the first approach with a discussion based on the underlying symmetry and topological principles, which are essential in designing a set of Josephson junctions to probe the monopole superconducting order.
@@ III.1 Microscopic derivation of Josephson coupling for monopole superconductors
We first extend the results from linear response theory in Sec. II to determine the first-order Josephson current for junctions that include monopole superconductors. Monopole pairing order is an exotic class of topologically obstructed pairing order that arises from the nontrivial Berry phase of Cooper pairs. This, for example, can arise when pairing occurs between two Fermi surfaces $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$ with different Chern numbers, $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ respectively, in doped time-reversal symmetry broken Weyl semimetals in three dimensions Li and Haldane (2018). In the weak-coupling regime, we describe this system using mean-field BdG Hamiltonian in Eq. (2), with its Nambu spinor modified to $\Psi(\mathbf{k})=(c_{1;\mathbf{k},\uparrow},c_{1;\mathbf{k},\downarrow},c_{2;-\mathbf{k},\uparrow}^{\dagger},c_{2;-\mathbf{k},\downarrow}^{\dagger})^{\mathrm{T}}$ to describe the inter-Fermi surface pairing $\Delta_{\mathrm{inter}}(\mathbf{k})$ between parity-related $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$, where $c_{i;\mathbf{k},\sigma}$ denotes the annihilation operator of an electron on $\mathrm{FS}_{i}$ ($i=1,2$) with momentum $\mathbf{k}$ and spin $\sigma=\uparrow,\downarrow$. Correspondingly, the BdG Hamiltonian kernel takes the following form,
@@
The inter-Fermi surface pairing, $\Delta_{\mathrm{inter};\sigma,\sigma^{\prime}}(\mathbf{k})\sim\langle c_{2;-\mathbf{k},\sigma^{\prime}}c_{1;\mathbf{k},\sigma}\rangle$ in the spin-$\uparrow$,$\downarrow$ , ↓ representation, describes the zero center-of-mass momentum pairing, since FS1 and FS2 are related by parity. In the weak-coupling regime, superconducting pairing occurs at low-energy close to the Fermi surfaces; hence, we express the pairing in the helical band representation as $\Delta_{\mathrm{inter};s,s^{\prime}}^{(b)}(\tilde{\mathbf{k}})\sim\langle\psi_{2;-\mathbf{\tilde{k}},s^{\prime}}\psi_{1;\mathbf{\tilde{k}},s}\rangle$, where $\psi_{1;\mathbf{\tilde{k}},s}=\sum_{\sigma=\uparrow,\downarrow}\chi_{1;s,\sigma}^{*}(\tilde{\mathbf{k}})c_{1;\tilde{\mathbf{k}}+\mathbf{K}_{0},\sigma}$ and $\psi_{2;\mathbf{\tilde{k}},s}=\sum_{\sigma=\uparrow,\downarrow}\chi_{2;s,\sigma}^{*}(\tilde{\mathbf{k}})c_{2;\tilde{\mathbf{k}}-\mathbf{K}_{0},\sigma}$ are annihilation operators of helical band eigenstates of $\mathcal{H}_{\mathrm{kin},1}$ and $\mathcal{H}_{\mathrm{kin},2}$ respectively. When the Chern numbers of two Fermi surfaces are different, the inter-Fermi surface pairing inherits the band topology of single-particle states at $\mathbf{k}$ and $-\mathbf{k}$ in a nontrivial way, and the Cooper pair scatters in the presence of a magnetic monopole in momentum space given by the pair monopole charge, $q_{p}=(\mathcal{C}_{1}-\mathcal{C}_{2})/2$. Consequently, this leads to a topologically obstructed superconducting pairing order that cannot be well defined over an entire Fermi surface. Instead, the pairing order is described by monopole harmonics Wu and Yang (1976, 1977) that all belong to the topological sector characterized by the pair monopole charge $q_{p}$ Li and Haldane (2018).
Upon projecting to the helical Fermi surfaces in the weak-coupling regime, the effective pairing in the helical band-diagonal Nambu basis $(\psi_{1;\tilde{\mathbf{k}},+},\psi^{\dagger}_{2;-\tilde{\mathbf{k}},+})^{\mathrm{T}}$ is
@@
To note, if each Fermi surface of the doped Weyl system has additional symmetry with respect to its enclosed Weyl point, intra-Fermi surface pairing is also possible, corresponding to nonzero center-of-mass momentum Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) pairing Fulde and Ferrell (1964); Larkin and Ovchinnikov (1964). As the pair monopole charge of the FFLO pairing is zero, the pairing can be fully gapped and compete with the inter-Fermi-surface pairing Li and Haldane (2018). However, without the additional symmetry of Fermi surfaces, the inter-Fermi-surface monopole pairing is energetically preferable to the intra-Fermi-surface FFLO pairing. Hence, we only consider zero center-of-mass pairing in this work.
@@
$=\sqrt{2\pi}\left(\mathcal{Y}_{\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}}),-\mathcal{Y}_{\frac{1}{2};\frac{1}{2},+\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\right),$
and is related to states on $\mathrm{FS}_{2}$ by $\chi_{2,+}(\tilde{\mathbf{k}})=\sigma_{z}\chi_{1,+}(-\tilde{\mathbf{k}})$. Here, $\bm{\Omega}_{\tilde{{k}}}=(\theta_{\tilde{k}},\varphi_{\tilde{k}})$, with $\theta_{\tilde{k}}$ and $\varphi_{\tilde{k}}$ being the polar and azimuthal angles of the spherical Fermi surfaces. $\mathcal{Y}_{q;l,l_{z}}(\bm{\Omega}_{\tilde{k}})$ is the monopole harmonic function Wu and Yang (1976, 1977) (reviewed in Appendix C), now defined in momentum space. The monopole harmonic function is labeled by the monopole charge $q$, the eigenvalue of angular momentum $l$, and the eigenvalue of $z$-component angular momentum $l_{z}$. Above, $\lambda=\pm 1$ corresponds to two different choices of gauge. The upper (lower) sign denotes the choice where the state $\chi_{1,+}(\tilde{\mathbf{k}})$ is well-defined locally near the north (south) pole of $\mathrm{FS}_{1}$, while $\chi_{2,+}(\tilde{\mathbf{k}})$ is well-defined near the south (north) pole of $\mathrm{FS}_{2}$, with the Dirac strings located at the respective antipodal points of these two Fermi surfaces. The Chern numbers of $\chi_{1,+}$ on $\mathrm{FS}_{1}$ and $\chi_{2,+}$ on $\mathrm{FS}_{2}$ are $\mathcal{C}_{1}=-1$ and $\mathcal{C}_{2}=+1$ respectively. The unitary transformations that diagonalize the Weyl band Hamiltonians $\mathcal{H}_{\mathrm{kin}}^{(\nu=\pm 1)}$ in Eq. (19) take the form
@@
where $\mathcal{Y}^{*}_{q;l,l_{z}}=(-1)^{q+l_{z}}\mathcal{Y}_{-q;l,-l_{z}}$.
@@
We now derive the form factor in the first-order Josephson coupling between a spin singlet superconductor with a spin-independent band Hamiltonian at the left side and the $q_{p}=-1$ monopole superconductor at the right side, where there is only bare spin-independent tunneling at the barrier. In the band-projected basis, the monopole pairing order is described by an effective $d$-vector shown in Eq. (17). The form factor hence takes the form
@@
$T^{(b)}_{0}(\pm\mathbf{k},\pm\mathbf{k}^{\prime})$ $=T_{0}(\pm\mathbf{k},\pm\mathbf{k}^{\prime})[u_{R,0}^{(\pm\mathbf{K}_{0})}(\pm\tilde{\mathbf{k}})]^{*},$ (26a)
@@
Here, $(u_{R,0}^{(\pm\mathbf{K}_{0})}(\tilde{\mathbf{k}}),\mathbf{u}_{R}^{(\pm\mathbf{K}_{0})}(\tilde{\mathbf{k}}))$ are the complex four-vectors defined in Eq. (22), related to the unitary transformation of the Weyl band Hamiltonian. As the spinful bands of the $s$-wave superconductor are decoupled, we have $u_{L,0}=1$ and $\mathbf{u}_{L}=0$.
Consequently, the form factor describing the first-order Josephson coupling between a $q_{p}=-1$ monopole superconductor and a singlet superconductor simplifies to
@@
×$\times{T_{0}(-\mathbf{k},-\mathbf{k}^{\prime})}{T_{0}(\mathbf{k},\mathbf{k}^{\prime})}$
×$\times\mathcal{Y}_{-\frac{1}{2};\frac{1}{2},-\frac{1}{2}}(\bm{\Omega}_{\tilde{k}})\mathcal{Y}_{-\frac{1}{2};\frac{1}{2},\frac{1}{2}}(\bm{\Omega}_{\tilde{k}}),$
where $\Delta_{L,0}(\mathbf{k}^{\prime})=d_{L,0}(\mathbf{k}^{\prime})$, $\Delta_{R,\mathrm{MSC}}^{{(q_{p}=-1)}}(\tilde{\mathbf{k}})$ is the band-projected effective monopole harmonic pairing order in Eq. (23) characterized by the pair monopole charge $q_{p}=-1$, and $\tilde{\mathbf{k}}=\mathbf{k}-\mathbf{K}_{0}$ is the wave vector with respect to the Weyl node. Despite the form factor expressed in Eq. (27) including multiple monopole harmonic functions, $\mathfrak{F}_{\mathrm{sing}-\mathrm{MSC}}$ is always non-singular, as it is related to the Josephson current, a physical observable. By employing the addition theorem of monopole harmonic functions Wu and Yang (1977) (see also Appendix C), which states that the product of two monopole harmonic functions with charges $q_{1}$ and $q_{2}$ can be expressed as a sum of monopole harmonic functions with monopole charge $q_{3}=q_{1}+q_{2}$, we find that the form factor in Eq. (27) has zero total monopole charge. In other words, it can be represented using standard spherical harmonic functions without singularities over the Fermi surface.
@@
$\mathfrak{F}_{\mathrm{trip}-\mathrm{MSC}}^{(q_{p}=-1)}(\tilde{\mathbf{k}},\mathbf{k}^{\prime})=-2\pi{T_{0}(-\mathbf{k},-\mathbf{k}^{\prime})}{T_{0}(\mathbf{k},\mathbf{k}^{\prime})}$
×$\times{\Delta_{R,\mathrm{MSC}}^{{(q_{p}=-1)}^{*}}(\tilde{\mathbf{k}})}[(d_{L,x}(\mathbf{k}^{\prime})-id_{L,y}(\mathbf{k}^{\prime}))\mathcal{Y}_{-\frac{1}{2};\frac{1}{2},-\frac{1}{2}}^{2}(\bm{\Omega}_{\tilde{k}})$
@@
It depends on, $\mathbf{d}_{L}(\mathbf{k}^{\prime})$, the $d$-vector which encodes the spin structure of the triplet pairing order at the left. In contrast, the form factor $\mathfrak{F}_{\mathrm{sing}-\mathrm{MSC}}^{(q_{p}=-1)}$ in Eq. (27) takes a simpler form, as the symmetry of the singlet pairing order at the left is primarily determined by its symmetry in momentum space. For the microscopic models studied later in Secs. IV and VII, to identify the momentum space topology of the monopole pairing order, we will focus on junctions with a spin singlet superconductor for simplicity.